# Approximating the coefficients of the Bessel functions

### v3

@@ 1.1. Main result
$F_{N}^{D}(x_{1},\ldots,x_{N})$ $=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$
@@
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)>1$.
@@
    For all $\nu\in\Gamma_{\text{even}}$,
@@
Then, we can deduce the convergence of the sequence $\{\frac{a(N)}{\theta N}\}_{N\geq 1}$ in terms of moments. Conversely, we have that if the sequence $\{\frac{a(N)}{\theta N}\}_{N\geq 1}$ converges in terms of moments, then we can determine the asymptotics of the coefficients of the Bessel functions $J_{a(N)}^{A^{N-1}(\theta)}$ in sufficiently small neighborhoods of the origin. We prove a generalized version of this result in Corollary˜10.18.
@@
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  2. (a”)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{A}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{A}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  4. (c”)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{BC}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{BC}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  6. (e”)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{D}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{D}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  8. (f”)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}((\epsilon+(1,\ldots,1))!)^{-1}[1]\partial^{\epsilon+(1,\ldots,1)}F_{N}^{D}}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}((\epsilon+(1,\ldots,1))!)^{-1}[1]\partial^{\epsilon+(1,\ldots,1)}F_{N}^{D}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
Then, (a’) and (a”), (c’) and (c”), (e’) and (e”), and (f’) and (f”) are equivalent to (a), (c), (e), and (f), respectively.
@@
The equivalences involving (a’), (c’), (e’), and (f’) are straightforward by the definition of the coefficients $c_{\epsilon\lambda}$. The equivalences involving (a”), (c”), (e”), and (f”) follow from $m_{\epsilon}^{A}(N)=(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{A}$, $m_{\epsilon}^{BC}(N)=(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{BC}$, $m_{\epsilon}^{D}(N)=(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{D}$, and $l_{\epsilon}^{D}(N)=((\epsilon+(1,\ldots,1))!)^{-1}[1]\partial^{\epsilon+(1,\ldots,1)}F_{N}^{D}$. ∎
@@ 1.4. Weak convergence to the free convolution
Assume that $\theta_{0},\theta_{1}\geq 0$ for all $N\geq 1$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\\ \frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$. Let $\mu_{a}$ and $\mu_{b}$ be compactly supported Borel probability measures over $\mathbb{R}$. Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\\ \delta\left(\frac{b(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
@@
Furthermore, we discuss generalizations of Corollaries˜1.7 and 1.8 in Subsection˜10.5 when $a(N)$ and $b(N)$ are randomly distributed based on exponentially decaying measures rather than deterministic. In particular, we use this framework to analyze the Dyson Brownian motion with random initial values in Example˜10.21. We establish that if the average of the scaled empirical distribution of the initial values converge in terms of moments to a Borel probability measure $\eta$ over $\mathbb{R}$, then the average of the scaled empirical distribution after $\alpha N$ time steps converges in terms of moments to a measure determined by $\eta$ and $\alpha$.
@@ 2.1. Dunkl operators
    If $\text{Re}(\theta(r))\geq 0$ for all $r\in\mathcal{R}$, then $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\max_{h\in H(\mathcal{R})}\text{Re}(\\ \left\langle ha,x\right\rangle))$.
@@ 8.1. Orthogonality results
  • The function $w:\mathbb{R}^{N}\rightarrow\mathbb{R}$ is defined as $w(x)\triangleq\frac{e^{-\left\lVert x\right\rVert_{2}^{2}}}{(2\pi)^{\frac{N}{2}}}$.
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}|\theta N|=\infty$. Consider the following statements.
@@
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=d_{\lambda}\in\mathbb{C}$.
@@
    For all $\nu\in\Gamma_{\text{even}}$,
@@
See Theorem˜8.11 for the equivalence of (a) and (c). For the equivalence of (b) and (d), we can use Theorem˜8.16 and follow the method of the proof of Theorem˜6.10, after dividing by the nonzero quantity $\prod_{j=1}^{N}(1+2(j-1)\theta)$. ∎
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}|\theta N|=\infty$. Consider the following statements.
@@
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)>1$.
@@
    For all $\nu\in\Gamma_{\text{even}}$,
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$. Consider the following statements.
@@
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$.
@@
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N-k}(1+2(j-1)\theta)}=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{D;\,\text{odd}}_{\nu\lambda}(2c)[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right).$
@@
Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer, then (d) implies (b).
@@
For the implications involving (a) and (c), see Theorem˜8.13. Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$. Using Theorem˜8.23 gives that $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ is divisible by $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$. Then, after dividing by this nonzero quantity, we can use Theorem˜8.23 and the same analysis as in the proof of Theorem˜6.17 to deduce that (b) implies (d). Furthermore, we similarly have that if $2c$ is not a negative integer, then $\mathcal{W}^{D;\,\text{odd}}(2c)[k]$ is invertible for all $k\geq 1$, so (d) implies (b) if this is the case. ∎
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$. Consider the following statements.
@@
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}d_{\lambda}(N)=0$ if $\ell(\lambda)>1$.
@@
    For all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N-k}(1+2(j-1)\theta)}$
@@
Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer, then (d) implies (b).

### v4

@@ 1.1. Main result
$F_{N}^{D}(x_{1},\ldots,x_{N})$ $=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)++ed_{0}(N)+e\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$
@@
    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)>1$.
@@
    It is the case that
    $\lim_{N\rightarrow\infty}\frac{[1]\prod_{j=1}^{N}\mathcal{D}_{j}(D^{N}(\theta))F_{N}^{D}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1$
    and for all $\nu\in\Gamma_{\text{even}}$,
@@
Then, we can deduce the convergence of the sequence $\{\frac{a(N)}{\theta N}\}_{N\geq 1}$ in terms of moments. Conversely, we have that if the sequence $\{\frac{a(N)}{\theta N}\}_{N\geq 1}$ converges in terms of moments, then we can determine the asymptotics of the coefficients of the Bessel functions $J_{a(N)}^{A^{N-1}(\theta)}$ in sufficiently small neighborhoods of the origin. We prove a generalized version of this result in Corollary˜10.17.
@@
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
@@
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
Then, (a’), (c’), (e’), and (f’) are equivalent to (a), (c), (e), and (f), respectively.
@@
The equivalences are straightforward after using the definition of the coefficients $c_{\epsilon\lambda}$. ∎
@@ 1.4. Weak convergence to the free convolution
Assume that $\theta_{0},\theta_{1}\geq 0$ for all $N\geq 1$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}\\ =c\in\mathbb{C}$. Let $\mu_{a}$ and $\mu_{b}$ be compactly supported Borel probability measures over $\mathbb{R}$. Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
@@
Furthermore, we discuss generalizations of Corollaries˜1.7 and 1.8 in Subsection˜10.5 when $a(N)$ and $b(N)$ are randomly distributed based on exponentially decaying measures rather than deterministic. In particular, we use this framework to analyze the Dyson Brownian motion with random initial values in Example˜10.26. We establish that if the average of the scaled empirical distribution of the initial values converge in terms of moments to a Borel probability measure $\eta$ over $\mathbb{R}$, then the average of the scaled empirical distribution after $\alpha N$ time steps converges in terms of moments to a measure determined by $\eta$ and $\alpha$.
@@ 2.1. Dunkl operators
    If $\text{Re}(\theta(r))\geq 0$ for all $r\in\mathcal{R}$, then $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\max_{h\in H(\mathcal{R})}\\ \text{Re}(\left\langle ha,x\right\rangle))$.
@@ 8.1. Orthogonality results
  • The function $w:\mathbb{R}^{N}\rightarrow\mathbb{R}$ is defined as $w(x)\triangleq\frac{e^{-\frac{\left\lVert x\right\rVert_{2}^{2}}{2}}}{(2\pi)^{\frac{N}{2}}}$.
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+ed_{0}(N)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}|\theta N|=\infty$. Consider the following statements.
@@
    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=d_{\lambda}\in\mathbb{C}$.
@@
    It is the case that
    $\lim_{N\rightarrow\infty}\frac{[e,F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1.$
    Furthermore, for all $\nu\in\Gamma_{\text{even}}$,
@@
See Theorem˜8.11 for the equivalence of (a) and (c). For the equivalence of (b) and (d), we can use Theorem˜8.16 and follow the method of the proof of Theorem˜6.10, after dividing by the nonzero quantity $\prod_{j=1}^{N}(1+2(j-1)\theta)$. In particular, we have that $e,F_{N}]_{D^{N}(\theta)}=(d_{0}(N)+1)[e,e]_{D^{N}(\theta)}=(d_{0}(N)+1)\prod_{j=1}^{N}(1+2(j-1)\theta))$, so we can deduce that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}$ and
$\lim_{N\rightarrow\infty}\frac{[e,F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1$
are equivalent. ∎
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+ed_{0}(N)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}|\theta N|=\infty$. Consider the following statements.
@@
    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{d_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)>1$.
@@
    It is the case that
    $\lim_{N\rightarrow\infty}\frac{[e,F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1.$
    Furthermore, for all $\nu\in\Gamma_{\text{even}}$,
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+ed_{0}(N)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$. Consider the following statements.
@@
    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$.
@@
    It is the case that
    $\lim_{N\rightarrow\infty}\frac{[e,F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1$
    and for all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N-|\nu|}(1+2(j-1)\theta)}$
    $=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{D;\,\text{odd}}_{\nu\lambda}(2c)[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right).$
@@
Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer, then (d) implies (b).
@@
For the implications involving (a) and (c), see Theorem˜8.13. Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Using Theorem˜8.23 gives that $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ is divisible by $\prod_{j=1}^{N-k}(1+2(j-1)\theta)$. Then, after dividing by this nonzero quantity, we can use Theorem˜8.23 and the same analysis as in the proof of Theorem˜6.17 to deduce that (b) implies (d). Furthermore, we similarly have that if $2c$ is not a negative integer, then $\mathcal{W}^{D;\,\text{odd}}(2c)[k]$ is invertible for all $k\geq 1$, so (d) implies (b) if this is the case. ∎
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+ed_{0}(N)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$. Consider the following statements.
@@
    It is the case that $\lim_{N\rightarrow\infty}d_{0}(N)=d_{0}\in\mathbb{C}$ and for all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}d_{\lambda}(N)=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}d_{\lambda}(N)=0$ if $\ell(\lambda)>1$.
@@
    It is the case that
    $\lim_{N\rightarrow\infty}\frac{[e,F_{N}]_{D^{N}(\theta)}}{\prod_{j=1}^{N}(1+2(j-1)\theta)}=d_{0}+1$
    and for all $\nu\in\Gamma_{\text{even}}$,
    $\lim_{N\rightarrow\infty}\frac{[ep_{\nu},F_{N}]_{D^{N}(\theta)}}{N^{\ell(\nu)}\prod_{j=1}^{N-|\nu|}(1+2(j-1)\theta)}$
@@
Assume that if $N$ is sufficiently large, then $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$. Then, (b) implies (d) and if $2c$ is not a negative integer, then (d) implies (b).